How do I solve this system of equations
x + 2y + z = 135
x + y + z = 118
2x + 2y + 3z = 269

Answer by Alan3354(69443)

How do I solve this system of equations
x + 2y + z = 135
x + y + z = 118
2x + 2y + 3z = 269
===================
Subtract the 2nd eqn from the 1st --> y = 17
------
Multiply the 1st eqn by 2, then subtract the 3rd.
=============================
I would use determinants.